Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 217     Entry time: 01 Jun 2023, 23:27
Author: Romain Letrun 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 10.65 keV, 128 pulses at 1.13 MHz and 16 pusles at 141 kHz.
  • Overall stable delivery.

Optical laser delivery

  • 800 nm, 50 fs, 1-226 pulses at 1.13 MHz.
  • A few issues with minor impact on operation this week. Detail communicated to LAS.

Achievements/Observations

  • Continued characterisation of crystals (thermal measurement, interferometry, imaging) under different beam conditions ranging from direct to focused beam.
  • Spectral measurement in parallel to crystals characterisation.
  • Imaging of static samples with Bragg magnifier.

Issues

  • DAQ configuration failing multiple times. Repeating the same sequence of actions yield variable results.
    • Several times with unhelpful message "Failed to execute "assign" action".
    • On one occasion with the message "Assignment for data-source/group 'SPB_CFEL_JF1M/DET/JNGFR09:daqOutput'/'SPB_CFEL_JF1M_9' not specified." even though we used this group before without any issues.
  • JUNGFRAU 1M acquisition could not be stopped. Restarting control and receiver servers and associated devices resulted in the control device being stuck in INIT state. Detector had to be power cycled to solved the issue.
ELOG V3.1.4-7c3fd00